OpenStudy (anonymous):

Is this correct so far? 8^(-4x-5)=4^(2x-0 =ln 8^(-4x-5)= ln 4^(2x-0) =(-4x-5) * ln 8 = (2x-0) * ln 4 =(-4x-5) / (2x-0) = ln 4/ ln 8...

OpenStudy (solomonzelman):

\(\large\color{slate}{ 8^{-4x-5}=4^{2x-0} }\) is this the problem ?

OpenStudy (anonymous):

yes, but the second exponent should be 2x-1

OpenStudy (solomonzelman):

here, how I would start. \(\large\color{slate}{ \left( 8\right)^{-4x-5}=\left( 4\right)^{2x-1} }\) \(\large\color{slate}{ \left( 2^{3}\right)^{-4x-5}=\left( 2^{2}\right)^{2x-1} }\) \(\large\color{slate}{ \left( 2\right)^{3(-4x-5)}=\left( 2\right)^{2(2x-1)} }\)

OpenStudy (solomonzelman):

then you don't need to use any logarithms.

OpenStudy (solomonzelman):

(and of course, then equation the exponents and solve for x ... ) hope this is helpful

OpenStudy (anonymous):

as in now I should only need 3(-4x-5) = 2(2x-1) into -12x-15=4x-2 into x=13/3?

OpenStudy (anonymous):

x= -13/3

OpenStudy (solomonzelman):

-12x-15=4x-2 -15=16x-2 -13=16x x=-13/16

OpenStudy (solomonzelman):

like this

OpenStudy (anonymous):

sorry ....stupid mistake on my part. I greatly appreciate you taking the time to help me. Thank you